OpenStudy (anonymous):

Ari is taking a group of people to the movies. Each child costs $4 and each adult is $9. How much will he pay if the group consists of 3 adults and 4 children?

OpenStudy (anonymous):

A. $91 B. $52 C. $48 D. $43

OpenStudy (anonymous):

3(9)+4(4)=43$

OpenStudy (anonymous):

3*9+4*4=27+16=43

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D

OpenStudy (anonymous):

$43